A chemical engineer is studying the following reaction: H,(g)+12(g) → 2 HI(g) At the temperature the engineer picks, the equilibrium constant K, for this reaction is 0.11. The engineer charges (fills) three reaction vessels with hydrogen and iodine, and lets the reaction begin. He then measures the composition of the mixture inside each vessel from time to time. His first set of measurements are shown in the table below Predict the changes in the compositions the engineer should expect next time he measures the compositions reaction vessel compound pressure 4.81 atm ↑ increase 3.39 atm 0.38 atm ↑ increase .97 atm 2.55 atm | ︵) ↑ increase 2.04 atm 4.88 atm | ︵) ↑ increase 3.46 atm 。↑ increase 0.23 atm expected change in concentration ↓ decrease ↓ decrease ↓ decrease ↓ decrease ↓ decrease ↓ decrease ↓ decrease ↓ decrease ↓ decrease 。(no change) (no change) 。(no change) (no change) O (no change) (no change) O (no change) O (no change) (no change) ↑ increase HI ↑ increase 8 HI ↑ increase IH HI ↑ increase
Answer #1

A)
Equilibrium Quotient expression is
Qp = p(HI)^2/p(H2)*p(I2)
= (0.38^2)/(4.81^1)*(3.39^1)
=0.0089

comparing Q and k
since Q is lesser than K
reaction will move in forward direction

Answer:
H2: decrease
I2: decrees
HI: increase

B)
Equilibrium Quotient expression is
Qp = p(HI)^2/p(H2)*p(I2)
= (2.04^2)/(3.97^1)*(2.55^1)
=0.411

comparing Q and k
since Q is greater than K
reaction will move in backward direction

Answer:
H2: increase
I2: increase
HI: decrease

C)
Equilibrium Quotient expression is
Qp = p(HI)^2/p(H2)*p(I2)
= (0.23^2)/(4.88^1)*(3.46^1)
=0.0031

comparing Q and k
since Q is lesser than K
reaction will move in forward direction

Answer:
H2: decrease
I2: decrees
HI: increase